OpenStudy (anonymous):

Hello I have another word problem (yuckkkk) Please help. A group has equal number of male and female. If 12 more males join the group and 7 females leave the group, then there will be 2 times more males than females.

OpenStudy (anonymous):

Using simple linear equation with one unknown.

OpenStudy (ash2326):

Let initially the no. of males= no. of females be "x" If 12 more males join the group and 7 females leave the group, no. of males now= (x+12) no. of females now= (x-7) Do you get this part?

OpenStudy (anonymous):

yes I got that part

OpenStudy (ash2326):

then there will be 2 times more males than females. so x+12=2(x-7) Can you solve this for x?

OpenStudy (anonymous):

my first though was x+12+x-7=3x

OpenStudy (anonymous):

im obviously wrong lol i hate word probs

OpenStudy (anonymous):

oh ok I see, I can solve for x now I just needed the equation.... I got mislead by thinking about the group as a whole

OpenStudy (ash2326):

No problem :D

OpenStudy (anonymous):

thank you =)

OpenStudy (ash2326):

Yeah :D